Hi, I need help w 2 b & c, 3 and 4. Please help, thanks!
Quiz 2-3 Review View No Spacing Headng 1 Part I: Answer the short answer question in a clear, conclse, yet comprehensive manner. Graphs aned calculations are mandatory whenever appropriate. The domestic demand and supply conditions for writing paper in Belgium are given as follows: Demand: P- 700-20 Supply: P- 40 +(Qs/5) 1) What are the equilibrium price and quantity in autarky? 2) If the nation can trade freely with the rest of the world (ROW) at a price of $120, Does Belgium or ROW have comparative advantage? b. a. What are the domestic production and domestic consumption respectively? Would Belgium import or export and what is the amount? For Belgium, moving from autarky to free trade, what are the effects of trade on the economic well- being of consumers, producers, and the nation? Please quantify your answers. Create a graph and refer to your graph in your answer. Suppose the autarky price for ROW is $150. Graphically show the international market: world price, the amount of imports and exports, the gains for Belgium and ROW. In addition, who gains more? c. 3) 4)
Answer #1

1) For Belgium, Demand : P=700-2Qd or, Qd=350-P/2

and Supply : P=40+(Qs/5) or, Qs=5P-200

For equilibrium , Demand = Supply

or, 350-P/2 = 5P-200

or, 700-P=10P-400

or, 11P=1100

or, P=100

and Q= 5*100-200=300

2.a) When Belgium trades freely in the world market at a price of $120 , it will gain profit. As cost of production of writing paper ($100) is less than free trade price ($120), Belgium has comparative advantage in the production of writing paper.

b) Domestic production will be at price $120 i.e, Qs=5P-200 or, Q=5*120-200=400

Domestic consumption will be at $120 i.e, Qd=350-P/2 = 350-120/2 = 290.

c) Belgium will export writing paper . Amount to be imported = 400-290=110

3) In Belgium , due to shift from autarky to free trade, quantity produced in Belgium increased with a greater profit being enjoyed by domestic producers.However, due to free trade domestic consumers will incur loss due to higher prices. This causes decrease in domestic consumption.The result is shown in the diagram below.

Pice es 120 lob 21口300 чоо Q uanh

4) when autarky price for ROW is $150, reverse phenomenon is seen. When ROW goes for free trade, world price will be $120. In this case , producers of ROW incur loss while consumers of ROW gain profit. The case is shown in the diagram below.

Prica uce Rよ a s 150 12o 7 Quant ons. awtan o n